MARS DESERT RESEARCH STATION

EVA Report – February 10th

EVA Report:
SOL: 11
Date: 02/09/2017
Person filling out the report: Mamatha Maheshwarappa
Crew members involved in the EVA: Arpan Vasanth and Mamatha Maheshwarappa
EVA leader: Mamatha Maheshwarappa
Begin: 10.30 am
End: 02.00 pm
Type of EVA: Walking + Rover (Deimos)
Purpose: Collection of Soil Samples and photographs for outreach activities
UTM Coordinates: 12 S; Northing: 4247314; Easting: 520331 (For Drone experiments)
12 S; Northing: 4255749; Easting: 518259 Zone/Sector
Summary: We started the EVA at 1030 hrs by capturing photographs/videos for outreach activities using drone near the “Zubrin’s Head” (towards the highway). Drone was controlled to be always ahead of the rover to understand the terrain better. Also, we simulated the application of Drone in search of a crew member in case of emergency situations. The second part of the EVA was towards “Lith Canyon” to collect soil samples for characterizing human commensal bacteria. GPS coordinates were noted at each location for future reference. The EVA was concluded after returning to the hab noting down the hab’s operational parameters.
Kind regards,
Mamatha

Crew Photos – February 10th

04 Sneha set to water plants in the GreenHab
06 Lovely connection between Mars and the Sun
07 Fantastic landscape captured by Sai Arun
08 Beautiful site of MDRS captured by Sai Arun
03 Mamatha carrying out internal microbial sample collection
02 Saroj driving the Rover during the last EVA of Crew 174
01 Breathtaking Sunrise captured by Sneha Velayudhan
05 Proud Saroj (Crew Scientist) representing The University of Alabama in Huntsville

Operations Report – February 10th

*Operations Report*
SOL: 12
Name of person filing report: Arpan Vasanth
Non-nominal systems: 3 x wall outlets in the Engineering Bay and static
tank pipe leakage.
Generator — functioning
Solar — not checked
Diesel – 50%
Propane – 54 psi
Gasoline (5 Gallon containers for ATV) – all the containers are full
Water (trailer) – Tank #1 – <05%
– Tank #2 – <05%
Water (static) – 90%
Trailer to Static Pump used – Yes
Water (loft) – Static to Loft Pump used – Yes
Water Meter: full – 0120350
ATV’s Used: 1 x 300
Oil Checked: No
ATV # Fuel Used Gals: <1.5
Tires Status: nominal
# Hours Used Day: 0
Phobos used: No
Hours: start = 13.2, end = 14.0
Beginning charge: 100%
Ending charge: 79%
Currently charging: No
Deimos used: Yes
Hours: start = 15.0, end = 16.9
Beginning charge: 100%
Ending charge: 67%
Currently charging: Yes
HabCar used? No
Notes and Comments: – ATV 300 had issues with battery connections and broke
down twice. The crew members think that is due to lose wire connection from
batteries. We suggest CapCom to put this information across to Crew 175 to
avoid any issues during future EVAs.
Summary of Operations activities: –

Kind regards,
Mamatha

Crew Photos – February 9th

Trial on Astrophotography by Sai Arun Dharmik at MDRS.

 

Mamatha collecting soil samples at Lith Canyon.

 

Arpan’s drone experiment used to locate a crew member.

 

Arpan and Mamatha – Get, Set, Go (EVA spirit)!

 

Good try from Saroj to skip from sol summary report.

 

View of Mars on a bright afternoon.

Sol Summary – February 9th

Sol Summary
Sol: 11
Earth Date: 02/09/2017
Written by : Sai Arun Dharmik (Crew Geologist)

“A happy life consists not in the absence, but in the, mastery of hardships”
                                                                                                                  -Helen Keller
The day started with a loud scream in a shrill voice of Sneha (crew biologist). I woke up from a dream and was very anxious about what could have happened. Well, it was Saroj (crew scientist) playing a prank on Sneha, as he hid in the air lock. Though I couldn’t recollect what I was dreaming about, I started wondering what could be like to dream on Mars. Some events that crossed my mind were, successfully cultivating on Mars, a huge dust storm hitting us, a starry night with new constellations, a highly unlikely visit by Aliens ? It was time for our daily briefing session and everyone gathered around the table and the commander started briefing about the activities scheduled for the day. Towards the end of briefing, it was time to assign the responsibility of reports to be sent to our CapCom on Earth. For the second time during the mission, I took up the responsibility to prepare Sol summary of the day.

As we are coming to the end of our mission, we were recalling our mission planning, preparation to reach Mars (MDRS Hab station) and our journey so far. The crew is proud to have made it to the ‘Mars’ and perform all the experiments as planned. Tasty breakfast consisting ‘Poha’ and tea was prepared by Arpan and Mamatha and Sneha.

The first task for the team was to prepare Arpan and Mamatha for the EVA which began at 10:30 hrs. The first activity of the EVA was to check for the status of the hab support system and then visit ‘Lith Canyon’ and ‘Zubrin’s head’. One of the main objective for the EVA was to simulate the application of drone in search of a crew member during an emergency situations and to collect soil samples for characterizing human commensal bacterial. This EVA also turned out to be one of the most successful for the crew, with all the mission objectives achieved as per the plan. Once the EVA crew left the hab, it was time for Sneha to water her plants and feed them with the entitled dosage of Vitamin D in the greenhab. At the same time, I was in the science dome studying the rock samples under the microscope and taking the magnetic susceptibility measurements of the samples collected during the previous EVAs.

After the EVA, everyone gathered for the lunch. It was pinto beans’ curry and rice today, prepared by Sneha and Saroj. During the lunch, the EVA crew members described the scenic locations they had explored and how they completed all the tasked successfully. Saroj was excited to explore the ‘Lith Canyon’ region after getting a good review from the EVA crew and described his plans for the last EVA (tomorrow) to be a memorable one.
Towards the end of the day, the entire crew was eagerly awaiting to update the mission support team on the day’s activities.

EVA Report – February 9th

EVA Report:
SOL: 11
Date: 02/09/2017
Person filling out the report: Mamatha Maheshwarappa
Crew members involved in the EVA: Arpan Vasanth and Mamatha Maheshwarappa
EVA leader: Mamatha Maheshwarappa
Begin: 10.30 am
End: 02.00 pm
Type of EVA: Walking + Rover (Deimos)
Purpose: Collection of Soil Samples and photographs for outreach activities
UTM Coordinates: 12 S; Northing: 4247314; Easting: 520331 (For Drone experiments)
12 S; Northing: 4255749; Easting: 518259 Zone/Sector
Summary: We started the EVA at 1030 hrs by capturing photographs/videos for outreach activities using drone near the “Zubrin’s Head” (towards the highway). Drone was controlled to be always ahead of the rover to understand the terrain better. Also, we simulated the application of Drone in search of a crew member in case of emergency situations. The second part of the EVA was towards “Lith Canyon” to collect soil samples for characterizing human commensal bacteria. GPS coordinates were noted at each location for future reference. The EVA was concluded after returning to the hab noting down the hab’s operational parameters.